OpenStudy (anonymous):

3(x+3)^2(x-2)^-3-3(x+3)^3(x-2)^-4=

OpenStudy (anonymous):

\[\frac{3 (x+3)^2}{(x-2)^3}-\frac{3 (x+3)^3}{(x-2)^4}\text{=}-\frac{15 (x+3)^2}{(x-2)^4} \]